# namerefThe definition of Are magnetic fields in cosmic voids primordial?

### v1

@@ Abstract
The nature of magnetic fields in the voids of the large-scale structure of the Universe has been a multifaceted open puzzle for decades. On one hand, their origin is not clear with most of the magnetogenesis models using physics beyond the standard model in the early Universe, and on the other hand, their existence and potential role in explaining the spectra of TeV blazars have been intensely debated in the past decade. Here, we show that void fields must exist and can be generated using classical electrodynamics in the late Universe. Specifically, we use the dipole component of the galactic fields, decaying as , and show that this generates space-filling magnetic fields in voids with sufficient amplitude to explain the blazar spectra and the lack of GeV halos observed by Fermi-LAT. Our results also show that the void fields might be too strong even for Cherenkov Telescope Array to observe the GeV halos. While primordial magnetic fields could still have been generated through some non-standard physics in the early Universe, our results show that the need for primordial magnetic fields is not strengthened by the presence of void magnetic fields.
@@ I Introduction
The standard model of cosmology (CDM) has been very successful in explaining a wide range of observations of the Universe and its history, including Hubble’s law of the expansion of space [1], numerous features of the cosmic microwave background [2], and the large-scale structure (LSS) of the Universe. However, there are still many open questions related to, for example, the inflationary stage at the beginning [3, 4], cosmological phase transitions that may have occurred later [5, 6], the asymmetry between matter and antimatter [7], nature of dark matter and dark energy [8], relative abundance of light elements [9], Hubble tension [10], tension [11], galaxy formation [12], and galactic magnetism [13, 14].
@@
It has been shown that magnetic fields can have an impact on several of the aforementioned phenomena and can, thus, help shape the Universe significantly (for reviews, see Refs. [15, 16, 17]). Some of the leading theories of inflation [e.g. 18] predict the production of magnetic fields that can survive until today in the voids of LSS [19, 20]. These primordial fields would also impact the spectrum of the relic gravitational waves [21] from inflation [22] as well as from (first-order) cosmological phase transitions [23]. If primordial magnetic fields are helical as predicted by some inflationary magnetogenesis models, they can even help explain the asymmetry between matter (baryons) and antimatter (antibaryons) through the conservation of the Chern-Simons number [24, 25]. If present, primordial magnetic fields have also been shown to play a role in many important tensions in modern cosmology, like the Hubble tension and the tension [26, 27, 28], and the lithium abundance problem [29]. Beyond their direct effect on the physics of the early Universe, primordial magnetic fields can influence how cosmic structures from in the post-recombination Universe. Through additional density fluctuations in the primordial plasma, they modify the seeds of cosmic structure formation, which affects galaxy formation [30] and the reionization epoch in the late Universe, and therefore the global 21cm signal [31, 32, 33].
@@
Primordial magnetic fields are important not only for astrophysics, but they are also relevant to fundamental physics at the microscopic levels. One of our leading ways to probe fundamental high-energy physics is via particle accelerators, which can reach energies up to TeV scales. However, the standard model of the particle physics (SM) may well remain valid for many orders of magnitude above the current particle accelerators, which will not be accessible by accelerators in the foreseeable future [34]. Consequently, exploring the physics beyond the SM (BSM) remains an uncertain task with accelerator physics. On the other hand, the early Universe can help us get insight into some of these questions, since the energy was much higher than accelerator energies (maybe even up to the Planck mass TeV). Since all the magnetogenesis models proposed in the literature employ BSM physics, a detection of these primordial fields would open a direct window into the extension of the SM.
@@
With such high significance of primordial magnetic fields for addressing critical questions studied across different communities, there are several ongoing efforts to detect them (for a review, see [35]). However, the only place in the present-day Universe where these primordial fields can be expected to be in their original form are in the voids of the LSS. In the other regions of the Universe, the signature of the primordial magnetic field is most likely contaminated by astrophysical processes in the late Universe. One of the main such processes is the dynamo mechanism, which amplifies magnetic fields exponentially by randomly stretching, twisting and folding magnetic field lines through turbulent plasma motion [36, 37, 38]. This has also been demonstrated in cosmological simulations [39, 40, 41]. These highly non-linear plasma processes make it practically impossible to deduce the initial conditions of the magnetic fields before the emergence of cosmic structures, leaving the void fields as the only potential relic of the primordial field.
@@
While theorized for decades, the first direct evidence for the existence of void fields was reported in 2010 [42, 43], providing a proof of concept of the viability of primordial fields as a probe of fundamental physics. There, the authors used the spectra of TeV blazars to deduce the lower limit of the field strength of void magnetic fields of G for coherence length scales above 1 Mpc. The high volume-filling factor required of void fields for these observations makes any astrophysical process disfavored over a primordial origin, which produces large scale fields throughout the Universe and, hence, also in voids [15, 16, 17]. However, other plasma processes have also been proposed to explain the observed TeV blazar spectra [44, 45]. Given the extreme range of density and energy scales involved, both analytical and numerical studies have not yet provided a conclusive resolution, and the existence of void fields remains an open problem. A clear proof of the existence of void fields would be the observation of GeV halos around TeV blazars, which would allow for a direct measurement of the void field strength.
@@
In this paper, we propose that void magnetic fields need to exist, resulting solely from classical electrodynamics in the late Universe. Specifically, we show that the dipole component of the galactic magnetic fields can produce space-filling void fields of the magnitude within the range suggested by the blazar spectra.
II Simple analytical estimate
@@
Let us consider spiral galaxies, like our Milky Way, surrounding a void of the LSS. These galaxies will have a magnetic field with a dipole component that decays as . The sum of all of the dipole fields of the galaxies will result in an extragalactic magnetic field (EGMF) with 100% volume-filling factor. We first estimate the root mean square (RMS) of this EGMF at the center of a large void, where it will be the weakest. In the next section, we use numerical tools to calculate the characteristic fields in a void and the associated power spectrum.
@@
Let us consider a large void. For our order of magnitude analysis in this section, we assume a spherical void of the radius Mpc. A void is generally surrounded by filaments and nodes, but for this crude estimate, let us consider galaxies to be distributed uniformly over the shell of the 20 Mpc sphere. A more realistic distribution of galaxies should provide approx. corrections and should only strengthen the fields at the center since the fields will cancel out most effectively for a uniform distribution. For a typical intergalactic distance of Mpc, we can easily obtain the number of galaxies surrounding the void
@@
where is the fraction of all galaxies that have appreciable magnetic dipoles. For now, we can crudely take all the dipoles to have a similar strength of G with . However, this could later be changed to incorporate some distribution of the dipole strength across galaxies. The dipole of each galaxy decays as to contribute from one galaxy at the center of the void, given by
@@
Here, is the distance from the center of the galaxy at which G. According to Refs. [46, 47, 48], the out-of-plane component of the galactic magnetic field remains of the order of G up to the distance of kpc, so we set kpc. We then have galaxy dipoles each contributing a magnetic field of the strength at the center of the void. In this simple estimate, we neglect possible correlations of the galactic dipole fields and assume a random orientation that is uniformly distributed across the solid angle. We can use the Box-Muller transformation [49] to find that each component of will follow a Gaussian distribution normalized to a fixed radius. Upon adding of , each component of the resulting field will be a normally distributed variable. The magnitude of this total magnetic field at the center will have the Maxwell-Boltzmann distribution with the RMS given by . We have also confirmed this numerically with Mathematica [50]. Combining all of this, we obtain a rough estimate for the resulting field
@@
This is an estimate for the RMS value of the magnetic field at the center of the void. Note that the RMS field (3c) will be minimal at the center of the void. In the next section, we calculate the RMS value of the magnetic field numerically in the entire void and estimate its power spectrum.
III Numerical analysis
@@
In order to calculate the magnetic field in the whole void numerically, we start with the relatively simple configuration assumed in the previous section, i.e., of a totally empty void with galaxies uniformly distributed over its surface. For simplicity, we consider a cubic void rather than a spherical one in the numerical analysis, with the assumption that the void is sufficiently large that its shape does not impact the magnetic fields appreciably in most of the void volume.
The void is modeled as a cube with galaxies spread on the surface of the cube as point particles. The number of galaxies is taken to be 504, which is approximately the number of galaxies expected around a spherical void with 20 Mpc, Mpc and . Each galaxy is modeled as a randomly oriented dipole with a constant strength G at kpc.
@@
turns out to be that of a white noise, , which is expected given the lack of any correlations imposed on the system. The characteristic magnetic field at the scale of 1 Mpc, defined as
for , is around G, about two and a half orders of magnitude stronger than the rough estimate of the field at the center of the void (3c). The characteristic field across length scales is shown in Fig. 1. Also, this enhancement is not surprising as Eq. 3c gives the minimal field in the void, while Eq. 6 gives the mean field on the scale of 1 Mpc.
IV Conclusion
@@
With our brief estimate, we have shown that it is indeed possible to generate void magnetic fields strong enough to explain the observed blazar spectra without resorting to physics beyond the SM. We use the dipole component of galactic fields to show that they should pervade throughout the voids, generating fields well above the lower bounds set by the blazar spectra.
While magnetic fields could still have been generated in the early Universe (possibly through non-standard physics), our results show that the need for primordial magnetic fields is not strengthened by the presence of void magnetic fields.

### v2

# Are magnetic fields in cosmic voids primordial?

@@ Abstract
The nature of magnetic fields in the voids of the large-scale structure of the Universe has been a multifaceted open puzzle for decades. On one hand, their origin is not clear with most of the magnetogenesis models using physics beyond the standard model in the early Universe, and on the other hand, their existence and potential role in explaining the spectra of TeV blazars have been intensely debated in the past decade. Here, we propose a mechanism, within classical electrodynamics, that could fill the voids with late-Universe fields and, under certain conditions, dispel the need for primordial fields altogether to explain the void fields. Specifically, we use the dipole component of the galactic fields to generate space-filling magnetic fields in voids with white-noise spectrum and sufficient amplitude to explain the lack of GeV halos around TeV blazars observed by Fermi-LAT. A definitive test for such fields in the voids will be the white-noise spectral shape, which will constrain possible plasma processes in the voids to the ones that allow for the propagation of these dipole fields into the voids.
@@ I Introduction
The standard model of cosmology (CDM) has been very successful in explaining a wide range of observations of the Universe and its history, including Hubble’s law of the expansion of space [1], numerous features of the cosmic microwave background [2], and the large-scale structure (LSS) of the Universe. However, there are still many open questions related to, for example, the inflationary stage at the beginning [3, 4], cosmological phase transitions that may have occurred later [5, 6], the asymmetry between matter and antimatter [7], the nature of dark matter and dark energy [8], the relative abundance of light elements [9], the Hubble tension [10], the tension [11], galaxy formation [12], and galactic magnetism [13, 14].
@@
It has been shown that magnetic fields can have an impact on several of the aforementioned phenomena and can, thus, help shape the Universe significantly (for reviews, see Refs. [15, 16, 17]). Some of the leading theories of inflation [see e.g. 18] predict the production of magnetic fields that can survive until today in the voids of LSS [19, 20]. These primordial fields would also impact the spectrum of relic gravitational waves [21] from inflation [22] as well as from (first-order) cosmological phase transitions [23]. If primordial magnetic fields are helical as predicted by some inflationary magnetogenesis models, they can even help explain the asymmetry between matter (baryons) and antimatter (antibaryons) through the conservation of the Chern-Simons number [24, 25, 26]. If present, primordial magnetic fields have also been shown to play a role in many important tensions in modern cosmology, like the Hubble tension and the tension [27, 28], and the lithium abundance problem [29]. Beyond their direct effect on the physics of the early Universe, primordial magnetic fields can influence how cosmic structures form in the post-recombination Universe. Through additional density fluctuations in the primordial plasma, they modify the seeds of cosmic structure formation, which affects galaxy formation [30] and the reionization epoch in the late Universe, and therefore the global 21cm signal [31, 32, 33].
@@
Beyond their significance in astrophysics, primordial magnetic fields also have implications for fundamental physics at microscopic scales. One of our leading ways to probe fundamental high-energy physics is via particle accelerators, which currently can reach energies up to TeV scales. However, the standard model of the particle physics (SM) may well remain valid for many orders of magnitude above the current particle accelerators, which will not be accessible by accelerators in the foreseeable future [34]. Consequently, exploring the physics beyond the SM (BSM) remains an uncertain task with accelerator physics. On the other hand, the early Universe can help us get insight into some of these questions, since the energy was much higher than accelerator energies (maybe even up to the Planck mass TeV). Since all the magnetogenesis models proposed in the literature employ BSM physics, a detection of these primordial fields would open a direct window into extensions of the SM.
@@
The high significance of primordial magnetic fields for addressing critical questions studied across different communities, prompts several ongoing efforts to detect them (for a review, see [35]). However, the only place in the present-day Universe where these primordial fields can be expected to be in their original form are in the voids of the LSS. In the other regions of the Universe, the signature of the primordial magnetic field is most likely contaminated by astrophysical processes in the late Universe. One of the main such processes is the dynamo mechanism, which amplifies magnetic fields exponentially by randomly stretching, twisting and folding magnetic field lines through turbulent plasma motion [36, 37, 38]. These highly non-linear plasma processes make it practically impossible to deduce the initial conditions of the magnetic fields before the emergence of cosmic structures, leaving the void fields as the only potential relic of the primordial field.
@@
While theorized for decades, the first evidence for the existence of void fields was reported in 2010 [39, 40], providing a proof of concept of the viability of primordial fields as a probe of fundamental physics. There, the authors used the spectra of TeV blazars to deduce the lower limit of the magnetic field strength in voids of G for coherence length scales above 1 Mpc. While the high volume-filling factor suggested by these observations favor a primordial origin, astrophysical processes in the late Universe like plasma instabilities [41] and galactic winds [42] have also been proposed to explain blazar observations and void magnetic fields respectively. Given the extreme range of density and energy scales involved, both analytical and numerical studies have not yet provided a conclusive resolution, and the existence and the nature of void fields remains an open problem. A clear proof of the existence of void fields would be the observation of GeV halos around TeV blazars, which would allow for a direct measurement of the void field strength, and simultanously rule out alternative interpretations of the missing GeV observations [43]. Recently, an observation has been reported in [44] that would suggest a strong likelihood of the involvement of galactic winds or other late Universe processes. Whether these processes pollute primordial fields or whether they act as the dominant magnetogenesis source, more analysis is still required to elucidate the nature of the cosmological magnetic fields.
@@
In this article, we propose a mechanism, within classical electrodynamics, that could pollute the primordial magnetic fields in the voids with late-Universe fields of the same order of magnitude or higher, and, under certain conditions, could dispel the need for primordial fields altogether to explain the void fields. Specifically, we show that the dipole component of the galactic magnetic fields can produce space-filling void fields of a magnitude within the range suggested by the blazar spectra.
II Set-up and assumptions
Let us consider spiral galaxies, like our Milky Way, surrounding a void of the LSS. These galaxies are likely to have a magnetic field with a dipole component that decays as . For the results obtained here, we assume that a small fraction of the total energy of magnetic fields is in the dipole component. This is motivated by models of the Milky Way magnetic field [45, 46], as well as simulations of galactic magnetic fields that show a significant dipole component, especially in the halo where it is difficult to observe such fields [47]. Specifically, models of the field indicate that an dynamo could be present on galactic scales, which generates an antisymmetric field in the halo with a dipole component [48]. We assume that the fraction of magnetic energy in the dipole fields even though, as we will see later, a much smaller is sufficient as well.
The other underlying assumption concerns the propagation of the magnetic fields into the voids. Within ideal magnetohydrodynamics (MHD), the flux passing through a plasma can be modified only through the plasma motion at the boundary because of the frozen-in theorem. So, the propagation of the dipole fields into the voids requires either plasma motion or effects beyond ideal MHD around the time when the galactic dipoles are forming. The required plasma motion in voids could result not only from galactic winds but also from structure formation [49]. Furthermore, if the voids have lower baryon to dark matter ratio than average, the screening of the fields by the baryonic plasma becomes weaker, enhancing the propagation of the fields. Also, if structure formation sets in at , and dipoles form before reionisation, matter is still neutral and cannot be treated within the MHD approximation. The presence of turbulence and primordial magnetic fields is another scenario in which the dipole fields can propagate into the voids through non ideal-MHD effects like reconnection. The exact details of the propagation depends on the specific velocity dispersion and on the magnetic fields already present in the plasma. A definitive physical evidence of these dipole fields would be the observation of the spectral shape reported here, which requires multiple observations like the one reported in [44]. Such observations will, then, constrain the plasma processes possible in the cosmic voids and whether they are consistent with the propagation of the dipole fields into the voids.
If the above conditions are met, the dipole fields act like a static background magnetic field maintained by the dipoles with consistent plasma boundary conditions. They need to be accounted for before using void fields as a proxy for the primordial fields, not unlike the CMB foregrounds. In the case of galactic winds, these dipole fields are additional physics that need to be properly accounted. The sum of all of the dipole fields of the galaxies will result in an extragalactic magnetic field (EGMF) with 100% volume-filling factor.
@@ II Set-up and assumptions
Next, we first estimate the root mean square (RMS) of this EGMF at the center of a large void, where it will be the weakest. Then, in Sec. IV, we use numerical tools to calculate the characteristic fields in a void and the associated power spectrum.
III Analytical estimate at the center of a void
@@
Let us evaluate the resulting magnetic fields at the center of a void, where it will be the weakest. For this, let us consider a large void. For our order-of-magnitude estimate in this section, we assume a spherical void of the radius Mpc. A void is generally surrounded by filaments and nodes, but for this crude estimate, let us consider galaxies to be distributed uniformly over the shell of the 20 Mpc sphere. A more realistic distribution of galaxies would provide approximately corrections and should only strengthen the fields at the center since the fields will cancel out most effectively for a uniform distribution. For a typical galaxy density, Mpc3 of galaxies comparable to the Milky Way, intergalactic distance Mpc, the number of galaxies surrounding the void would be
@@
For now, we can crudely take all the galaxies to have a magnetic field of a similar strength of with a typical fraction of the energy in the dipole component given by , yielding the dipole field to be . In the next section, we will conduct numerical analysis using a lognormal distribution for the galactic field strength. The dipole of each galaxy decays as to contribute from one galaxy at the center of the void, given by
@@
Here, is the distance from the center of the galaxy at which . According to Refs. [46, 50, 51], the out-of-plane component of the galactic magnetic field remains of the order of G up to the distance of kpc, so we set kpc. We then have galaxy dipoles each contributing a magnetic field of the strength at the center of the void. In this simple estimate, we neglect possible correlations of the galactic dipole fields and assume a random orientation that is uniformly distributed across the solid angle. We can use the Box-Muller transformation [52] to find that each component of will follow a Gaussian distribution normalized to a fixed radius. Upon adding of , each component of the resulting field will be a normally distributed variable. The magnitude of this total magnetic field at the center will have the Maxwell-Boltzmann distribution with the RMS given by . Combining all of this, we obtain a rough estimate for the resulting field
@@
This is an estimate for the RMS value of the magnetic field at the center of the void. Note that the RMS field (3) will be minimal at the center of the void. While the amplitude of the spectrum will be larger in the rest of the void, its dependence on various parameters of the model would still satisfy Eq. 3. In the next section, we calculate the RMS value of the magnetic field numerically in the entire void and estimate its power spectrum.
IV Spectrum of the magnetic field in the whole void
@@
In order to calculate the RMS magnetic field from these dipoles and its coherence scale in the whole void numerically, we start with the relatively simple configuration assumed in the previous section, i.e., of a totally empty void with galaxies uniformly distributed over its surface. For simplicity, we consider a cubic void rather than a spherical one in the numerical analysis, with the assumption that the void is sufficiently large that its shape does not impact the magnetic fields appreciably in most of the void volume. The void is modeled as a cube with galaxies spread on the surface of the cube as point particles. The number of galaxies is taken to be 204, which is approximately the number of galaxies expected around a spherical void with 20 Mpc and .
Each galaxy is assumed to have a magnetic field of random strength with a lognormal probability distribution centering around
As expected, log of the expected strength of the dipole field depends on the variance of the normal distribution on the right hand side here, as explained in Appendix A. Each of these galactic fields are assumed to have a dipole component that is randomly oriented with a uniform distribution over the surface of a sphere and is of the strength at a distance from the center of the galaxy. We considered three different sets of values for these parameters, tabulated in Table 1.
| Parameter | Setup A | Setup B | Setup C |
| 7 | 2 | 10 | |
| 1 | 0 | 2 | |
| 3 | 1 | 10 |
@@
is calculated from the average of the spectrum over multiple ensembles. Details about the ensemble averaging and the convergence of the results are reported in Appendix B.
@@
The resultant spectrum turns out to be that of a white noise, , which is expected given the lack of any correlations imposed on the system. The characteristic magnetic field for a given coherence length scale is then defined as
This characteristic field across various length scales is shown in Fig. 1 as purple lines, with the solid line denoting the results for setup and the dashed lines denoting setups and . For setup , the characteristic field at the scale of 1 Mpc is around , about three and a half orders of magnitude stronger than the rough estimate of the field at the center of the void (3). This enhancement is not surprising as Eq. 3 gives the minimal field in the void, while Eq. 7 gives the RMS field on the scale of 1 Mpc. We expect the length scales to be certainly greater than that of a typical galaxy and perhaps a fraction of the distance between dipoles. We have cut off the length scales at because our simple model presented here cannot account for the physics on smaller scales. The lower hatched region (blue with vertical hatching) in Fig. 1 represents the region excluded by the earlier observations of TeV blazars, which provides a lower limit for void fields [40]. It has been suggested in [53] that, for smaller scales, the lower limit could have a different slope or shift to the right depending on the spectral index of the magnetic fields present. The diagonal red line denotes the size of the largest processed eddy at recombination, and, thus, represents the approximate region where primordial fields could be expected [54, 55]. Finally, the grey dashed line represents the theorized sensitivity below which Cherenkov Telescope Array Observatory (CTAO) is expected to be able to detect the void magnetic fields [26].
As shown in Fig. 1, the slope of the spectrum of the magnetic fields remains that of a white noise for the different setups given in Table 1. Additional physical considerations like the intrinsic alignment of galaxies and alignment of galaxy spins will make the spectrum flatter; however, those effects are beyond the scope of this article and are deferred to future work. The absolute value of the field strength, on the other hand, changes between the setups. Setups and depict the range we expect the parameters to take. The dependence on individual parameters would be the same as given by Eq. 3 with two exceptions. First, the amplitude would change from to the scale-dependent value depicted by the solid purple line in Fig. 1. Second, the linear dependence on the galactic field in Eq. 3 would be replaced by a linear dependence on from the left-hand side of Eq. 4 plus an additional dependence on the from the right-hand side of Eq. 4. This additional dependence on is detailed in Appendix A.
The observation of void fields reported in [44] is approximately of the strength of , which would suggest around orders of magnitude weaker field than that is depicted by the solid purple line. This could be achieved by many different combinations of the various parameters described in Eq. 3 since this value falls well within the range depicted by setups and . More observations would be required to ascertain the spectral slope. Overall, the parameter region of the observation of [44] suggests late Universe processes at play, which would strengthen the likelihood of these dipole fields to be the feature being observed.
V Conclusion
@@
With our brief estimate, we have shown that it is indeed possible to generate void magnetic fields strong enough to explain the observed blazar spectra. We use the dipole component of galactic fields to show that they could pervade throughout the voids, generating fields well above the lower bounds set by the blazar spectra.
The assumption of a dipole component of the galactic magnetic field is motivated by the models of the field of the Milky Way, as well as simulations of the galactic halos, which show the possibility of an dynamo that produces dipole fields among other components. Furthermore, if these dipole fields could propagate into the voids through plasma motions (including galactic winds) and non-ideal MHD effects, then these fields could explain the void fields without requiring primordial fields and, consequently, physics beyond the SM. In case primordial magnetic fields are present, these dipole fields could alter their signature significantly, requiring a careful cleanup of this “foreground.”
The model assumed here (setup in Table 1) uses of the magnetic energy to be in the dipole component, which leads to around at the scale of . A much smaller fraction of the total magnetic energy in the dipole component would still be enough to provide strong enough fields to explain the blazar observations and the reported observation in [44]. The dependence of the strength of the field on the various parameters of the model is given by Eq. 3, and is denoted by purple lines in Fig. 1. If the magnetic fields are sufficiently strong, stronger than , they would also suppress some of the plasma instabilities. The shape of the spectrum is found to be that of a white noise. This spectral shape would be the definitive proof of these dipole fields in the voids, requiring multiple observations like that of [44].
The results discussed here can be extended to include a realistic distribution of the galaxies, which would strengthen the estimates mentioned here because of the few galaxies present in the voids, intrinsic alignment of the galaxies and correlations among the matter overdensities. Another interesting extension would be to study a more realistic model of the interaction of the dipole fields with the plasma in voids to understand the significance of primordial fields, galactic winds and structure formation on the propagation of the dipole fields in the voids.
Acknowledgements
We thank Sergio Martin Alvarez, Rainer Beck, Abhijit Bendre, Chiara Caprini, Glennys Farrar, Andrii Neronov, Pranjal Ralegankar, Dimitri Semikov, Günter Sigl and Kandu Subramanian for important comments on the first draft.
Appendix A Impact of model parameters
The expected strength of the dipole field in a void depends on the distribution assumed for the galactic fields in Eq. 4. While keeping other parameters as per setup of Table 1, we varied the variance of the normal distribution assumed on the right hand side of Eq. 4, and found that the log (to the base 10) of the strength of the void field varied as approximately , as shown in Fig. 2. Here, the square root of two appears because we normalized the strength with respect to the case of that is used in the analysis presented in the main text. While the general form of this dependence is expected as we considered a lognormal distribution for the galactic fields and the void fields depend linearly on the galactic fields, the specific value of is an empirical fit.
Appendix B Convergence of the ensemble average
To calculate the ensemble average, different realizations of a void (setup in Table 1) are considered, and the root mean square of the magnetic field is calculated from those realizations. Each realization of a void consists of a specific realization of 204 surrounding galactic dipoles modeled as stochastic variables. Each of these dipoles have a direction that is uniformly distributed over the surface of a sphere and a magnitude given by , where is assumed to have the distribution given in Eq. 4. We find that the ensemble average converges to a stable value with less than 10% scatter after only two ensembles, which demonstrates the robustness of the averaging over 204 galaxies. To test this, we consider a subset of n ensembles (where n varies from two to nineteen) and we calculate the mean of each of different subsets of n ensembles. Then, we calculate the deviation of these 500 or less means from the full mean of the twenty ensembles.(We considere root mean square of the deviation across wavenumbers.) The spread of this deviation for 95% of the ensembles is shown as the purple region in Fig. 3. The mean deviation for a subset of n ensembles is shown as the blue line. The vertical axis is normalized to the full mean of the twenty ensembles so the tick-labels represent relative deviation from the full mean.
Expectedly, the deviation drops with increasing number of ensembles in the subset, as shown in Fig. 3. If we just consider two ensembles, the mean deviation from the full mean of twenty ensembles is already below 10%, suggesting that twenty ensembles yield very close to the true mean of the system.